OpenStudy (anonymous):

Why does: (2/5)^-3 = 125/8? I got up to 8/125, but why is it 125/8?

OpenStudy (neer2890):

\[2^{-1}=\frac{ 1 }{ 2 }\] so,\[(\frac{ 2 }{ 5 })^{-3}=\frac{ 1 }{ (\frac{ 2 }{ 5 })^{3} }=\frac{ 1 }{ \frac{ 8 }{ 125 } }=\frac{ 125 }{ 8 }\]

OpenStudy (anonymous):

Thank you neer2890! :) Makes so much sense now.

OpenStudy (neer2890):

you're welcome...:)

OpenStudy (anonymous):

Sorry, before you go, could you please explain what happens after 1/(8/125)?

OpenStudy (neer2890):

ok.\[\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }=\frac{ a }{ b }*\frac{ d }{ c }\]

OpenStudy (neer2890):

when in division, we just reciprocal it and multiply with numerator

OpenStudy (anonymous):

Thank you! :)

OpenStudy (neer2890):

you're welcome